OpenStudy (anonymous):

What is the solution to the system of equations? 5x-3y=-9 2x-5y=4

OpenStudy (anonymous):

@One098

OpenStudy (anonymous):

Sorry if it takes me awhile to answer my internet is bad.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

First Step: We have to make one of the variables cancel out when we add the two equations together.

OpenStudy (anonymous):

We do this by multiplying the top equation by 5, and the the bottom equation by (-3). Then when we add the two together the y's will cancel out. The equations are now: \[25x-15y=45\] and \[-6x+15y=-12\] Then when we add the two together we have: \[19x=33\] Understand so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Step Two: Solve for x.

OpenStudy (anonymous):

divide 19 on both sides right

OpenStudy (anonymous):

\[x = \frac{ 33 }{ 19 }\]

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

So x = 1.74

OpenStudy (anonymous):

plug in x

OpenStudy (anonymous):

Step Three: Take that 'x' value we just found, and put it back into one of the two original equations, and solve for 'y'.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

That looks like this: \[2(1.74)-5y= 4\] \[3.48 - 5y = 4\] \[-5y = 0.52\] \[y = -0.104\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So these two equations cross at (1.74,-0.104)

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

its not letting me put it all in

OpenStudy (anonymous):

Yup, we can check by graphing the two individual equations. Go here: https://www.desmos.com/calculator and type in 5*x-3*y = 9 for the first equation and 2*x - 5*y = 4.

OpenStudy (anonymous):

its -3 -2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I didn't notice the (-) in the first equation sorry! :S